# Sequential Monte Carlo with Gaussian Mixture Approximation for Infinite-Dimensional Statistical Inverse Problems

### v3

@@ 5.4.1 Discussion of total variation distance
Notice that $\mu_{1}^{N}=L_{1}S^{N}P_{0}\mu_{0}$ and the aforementioned iteration will terminate at
$d(\mu_{1}^{N},\mu_{1})\leq 2\kappa_{1}^{-1}\kappa_{2}^{-1}\left[\frac{1}{\sqrt{N}}+d(\mu_{0},\mu_{0})\right]=2\kappa_{1}^{-1}\kappa_{2}^{-1}\frac{1}{\sqrt{N}}.$
After iterating, we deduce that
$d(\mu_{J}^{N},\mu_{J})$ $\leq\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\left(\frac{2}{\kappa_{1}\kappa_{2}}\right)^{j}.$
This completes the proof. ∎
@@
$\mu_{0}(du)Q(u,dv)=$ $\sum_{j=1}^{M}w_{j}\mu_{j}(du,dv),\quad\mu_{j}=\mathcal{N}\left(m_{j},\mathcal{V}_{j}\right),$ (5.7)
$\mu_{0}(dv)Q(v,du)=$ $\sum_{j=1}^{M}w_{j}\mu_{j}^{\prime}(du,dv),\quad\mu_{j}^{\prime}=\mathcal{N}\left(m_{j}^{\prime},\mathcal{V}_{j}^{\prime}\right),$ (5.8)
@@
Then it follows from Theorem 5.d in Nguyen2020approximation that the density function can be approximated by $g_{k}$, i.e.,
@@
Note that the transition kernel $Q$ is part of $P$. In fact we have
$P(u,dv)=Q(u,dv)a(u,v)+\delta_{u}(dv)\int(1-a(u,w))Q(u,dw),$ (5.29)
where
$a(u,v)=$ $\min\left\{1,\frac{\mu_{post}(dv)Q(v,du)}{Q(u,dv)\mu_{post}(du)}\right\}=\min\left\{1,\frac{\mu_{post}(dv)}{Q(dv)}\frac{Q(du)}{\mu_{post}(du)}\right\}.$
For simplicity we use $Q(dv)$ instead of $Q(u,dv)$:
$P(u,dv)=Q(dv)a(u,v)+\delta_{u}(dv)\int(1-a(u,w))Q(dw).$ (5.30)
Now we compute the total variation distance between $Q\mu^{N}$ and $P\mu^{N}$:
$\int_{U}f(u)Q\mu^{N}(du)-\int_{U}f(u)P\mu^{N}(du)$
$=$ $\int_{U}f(u)\int_{V}Q(du)\mu^{N}(dv)-\int_{U}f(u)\int_{V}P(v,du)\mu^{N}(dv)$
$=$ $\iint_{UV}f(u)\left[1-a(v,u)\right]\mu^{N}(dv)Q(du)-\int_{V}f(v)\mu^{N}(dv)\int_{W}[1-a(v,w)]Q(dw)$
$=$ $\iint f(u)\left[1-a(v,u)\right]\mu^{N}(dv)Q(du)-\iint f(v)[1-a(v,u)]Q(du)\mu^{N}(dv)$
$=$ $\iint[f(u)-f(v)][1-a(v,u)]\mu^{N}(dv)Q(du).$
If $a=1$, then $d_{\text{\tiny TV}}=0$ and the result is obvious; otherwise $\frac{\mu_{post}(du)}{Q(du)}\frac{Q(dv)}{\mu_{post}(dv)}<1$ and
$d(Q\mu^{N},P\mu^{N})=$ $\sup_{|f|_{\infty}\leq 1}\left|\iint[f(u)-f(v)][1-a(v,u)]\mu^{N}(dv)Q(du)\right|$
$\leq$ $\iint 2|1-a(v,u)|\left|\mu^{N}(dv)-\mu_{post}(dv)\right|Q(du)$
$+\sup_{|f|_{\infty}\leq 1}\left|\iint(f(u)-f(v))(1-a(v,u))\mu_{post}(dv)Q(du)\right|$
$\leq$ $2d(\mu^{N},\mu_{post})+\sup_{|f|_{\infty}\leq 1}\left|\iint(f(u)-f(v))(1-a(v,u))\mu_{post}(dv)Q(du)\right|$
where the boundedness of $a(u,v)$ is used. Finally, a straight computation complete the proof:
$\sup_{|f|_{\infty}\leq 1}\left|\iint[f(u)-f(v)]\left[1-\frac{\mu_{post}(du)}{Q(du)}\frac{Q(dv)}{\mu_{post}(dv)}\right]\mu_{post}(dv)Q(du)\right|$
$=$ $\sup_{|f|_{\infty}\leq 1}\left|\iint[f(u)-f(v)]\left[\mu_{post}(dv)Q(du)-\mu_{post}(du)Q(dv)\right]\right|$
$=$ $\sup_{|f|_{\infty}\leq 1}\left|\int fdQ-\int fd\mu_{post}-\int fd\mu_{post}+\int fdQ\right|$
$=$ $2d(Q,\mu_{post}).$

### v4

@@ 5.4.1 Discussion of total variation distance
$(\text{triangle inequality})\quad$ $\leq\frac{1}{\sqrt{N}}+2\kappa_{1}^{-1}\kappa_{2,j}^{-1}[d(S^{N}P_{j}\mu_{j}^{N},P_{j}\mu_{j}^{N})+d(P_{j}\mu_{j}^{N},\mu_{j})]$
[…]
Notice that $\mu_{1}^{N}=RL_{1}S^{N}P_{0}\mu_{0}$ and the aforementioned iteration will terminate at
$d(\mu_{1}^{N},\mu_{1})\leq\frac{1}{\sqrt{N}}+2\kappa_{1}^{-1}\kappa_{20}^{-1}\left[\frac{1}{\sqrt{N}}+d(\mu_{0},\mu_{0})\right]=\frac{1}{\sqrt{N}}\left(1+2\kappa_{1}^{-1}\kappa_{20}^{-1}\right).$
Iterating completes the proof, and the result is given by
$d(\mu_{J}^{N},\mu_{J})\leq\frac{1}{\sqrt{N}}\Biggl[1+2\sum_{j=1}^{J-1}\frac{2^{j}}{\kappa_{1}^{j}}\prod_{k=1}^{j}\frac{1}{\kappa_{2,J-k}}+\prod_{j=0}^{J-1}\frac{2}{\kappa_{1}\kappa_{2,j}}\Biggr].$
In the following, under an intuitive setting, we illustrate the usefulness of the obtained error bounds, thereby demonstrating the necessity of developing the multi-layer SMC algorithm. Specifically, we consider the inverse problem associated with Darcy flow. For fixed values of $\kappa_{1}$ and $\kappa_{2,j}$, the qualitative dependence of the derived error bounds on $J$ can be inferred. Specifically, we set $\kappa_{1}=2$ and $N=5\times 10^{6}$, and assign all $\kappa_{2,j}$ a common value, denoted by $\kappa_{2}^{c}$. The resulting upper bounds on the error for different numbers of layers are reported in Table 2. The appropriateness of the chosen parameter values is demonstrated in the following Remark.
Table 2: Dependence of the approximation error on $J$.
  Number of layers $J$ The common lower bound $\kappa_{2}^{c}$ Error bound $d(\mu_{J}^{N},\mu_{J})$
1 0.001 1.4156
10 0.6 0.9298
20 0.9 0.1941
30 0.99 0.0990
40 0.999 0.1154
50 0.9999 0.1418
Thus, under this setting our error bound shows that SMC with multiple layers is strictly more accurate than the trivial single-layer case. It can also be observed that the error initially decreases as the number of layers increases, but eventually begins to increase again, indicating the existence of an optimal number of layers that minimizes the error. This behavior can be explained by the fact that our error bounds account for resampling at each layer, which inevitably introduces an error of order $O(1/\sqrt{N})$ per layer. When the number of layers becomes sufficiently large, these layerwise errors accumulate, offsetting the benefits of adding more layers.
Remark: This phenomenon, where the SMC error first increases and then decreases with the number of layers $J$, underlies the choice of $\kappa_{2}^{c}=0.001,0.6,0.9,0.99,0.999,0.9999$ for $J=1,10,20,30,40,50$. Recall that the constant $\kappa_{1}^{-1}$ serves as an upper bound for $\exp\!\left(-\Phi_{j}(u)\right).$ Since the observed data are contaminated by noise, they often do not belong to the range of the forward operator. Consequently, the potential function $\Phi_{j}(u)=\frac{h_{j}}{2\sigma^{2}}\|\mathcal{F}(u)-d\|^{2}$ may admit a strictly positive lower. With respect to the distribution of the noise, this occurs with high probability. It is therefore possible to choose $\kappa_{1}=2$. Indeed, the larger the smoothness gap between the noise and the forward solution, the greater the constant $\kappa_{1}$. Recall that $\kappa_{2,j-1}$ is the lower bound of the integral $\int\exp\!\left(-h_{j}\Phi(u)\right)\,\mu_{j-1}(du).$ In the Darcy flow example, the evaluation of $\Phi(u)$ requires the numerical solution of a PDE, which renders an analytic computation of $\kappa_{2,j}$ intractable. Nevertheless, the qualitative behavior is evident: as the total number of layers $J$ increases, the step sizes $h_{j}$ decrease, and consequently $\kappa_{2,j}$ approaches unity.
@@
$\mu_{0}(du)Q(u,dv)=$ $\sum_{j=1}^{M}w_{j}\rho_{j}(du,dv),\quad\rho_{j}=\mathcal{N}\left(m_{j},\mathcal{V}_{j}\right),$ (5.6)
$\mu_{0}(dv)Q(v,du)=$ $\sum_{j=1}^{M}w_{j}\rho_{j}^{\prime}(du,dv),\quad\rho_{j}^{\prime}=\mathcal{N}\left(m_{j}^{\prime},\mathcal{V}_{j}^{\prime}\right),$ (5.7)
@@
Then it follows from Theorem 5.d in Nguyen2020 that the density function can be approximated by $g_{k}$, i.e.,
@@
The proof of the first inequality (2.41) is straightforward, i.e.,
$d(Q,P\mu_{j}^{N})\leq d(Q,\mu_{j})+d(P\mu_{j},P\mu_{j}^{N})\leq d(Q,\mu_{j})+d(\mu_{j},\mu_{j}^{N}).$
To prove the second inequality (2.42), let $X_{1},\dots,X_{N}$ be samples obtained by SMC and define the empirical measure
$\mu^{N}\mathrel{\mathop{\ordinarycolon}}=\frac{1}{N}\sum_{i=1}^{N}\delta_{X_{i}}.$
Let $\rho=\mathcal{N}(0,\mathcal{C}_{0})$, and define the Gaussian mixture
$Q=\mu^{N}*\mathcal{\rho}=\frac{1}{N}\sum_{i=1}^{N}\bigl(\delta_{X_{i}}*\mathcal{N}(0,\mathcal{C}_{0})\bigr)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{N}(X_{i},\mathcal{C}_{0}),$
where the convolution of two finite Borel measures is defined by
$(\nu_{1}*\nu_{2})(A)\mathrel{\mathop{\ordinarycolon}}=\int\nu_{1}(A-x)\,\nu_{2}(dx).$
Difference is estimated by
$d^{\prime}(Q,\mu)\leq$ $d^{\prime}(Q,\mu*\rho)+d^{\prime}(\mu*\rho,\mu).$
The first term is bounded by
$d^{\prime}(\mu^{N}*\rho,\mu*\rho)$ $=\sup_{|f|_{\infty}\leq 1,\|D^{2}f\|\leq 1}\sqrt{\mathbb{E}\Big[\int fd(\mu^{N}*\rho)-\int fd(\mu*\rho)\Big]^{2}}$
$=\sup_{|f|_{\infty}\leq 1,\|D^{2}f\|\leq 1}\sqrt{\mathbb{E}\Big[\int(f*\rho)d(\mu^{N}-\mu*\rho)\Big]^{2}}$
$\leq d^{\prime}(\mu^{N},\mu)$ (5.28)
where $\rho(-A)=\rho(A)$ for all $A\in\mathcal{B}(\mathcal{H})$, $|f*\rho(x)|=\left|\int f(x-y)\rho(dy)\right|\leq 1$, and $D^{2}(f*\rho)=D^{2}f*\rho$ are employed.
Using once more the identity $\rho(-A)=\rho(A)$ for every $A\in\mathcal{B}(\mathcal{H})$, the second term is bounded by
$d^{\prime}(\mu*\rho,\mu)=$ $\sup_{|f|_{\infty}\leq 1,\|D^{2}f\|\leq 1}\left|\int(f*\rho-f)\mu(dx)\right|$
$=$ $\sup_{|f|_{\infty}\leq 1,f\in C_{\infty}}\left|\int_{x\in\mathcal{H}}\left(\int_{y\in\mathcal{H}}f(x-y)\rho(dy)-f\right)\mu(dx)\right|$
$=$ $\sup_{|f|_{\infty}\leq 1,\|D^{2}f\|\leq 1}\left|\int_{x\in\mathcal{H}}\int_{y\in\mathcal{H}}\left(f(x+y)-f(x)\right)\rho(dy)\mu(dx)\right|.$
Using a Taylor expansion together with $\mathbb{E}_{Y\sim\rho}Y=0$, we obtain
$d^{\prime}(\mu*\rho,\mu)=$ $\sup_{|f|_{\infty}\leq 1,\|D^{2}f\|\leq 1}\left|\int_{x\in\mathcal{H}}\int_{y\in\mathcal{H}}\frac{1}{2}(D^{2}f(x+ty)y,y)\rho(dy)\mu(dx)\right|$
$\leq$ $\frac{1}{2}\int\|y\|^{2}\rho(dy)=\frac{1}{2}\text{Tr}(\mathcal{C}_{0}).$
We choose $\rho=\mathcal{N}(0,\mathcal{C}_{0})$ such that $\mathrm{Tr}(\mathcal{C}_{0})\leq 2\epsilon\,d(\mu^{N},\mu)$, which implies that
$d^{\prime}(\mu*\rho,\mu)\leq\epsilon d^{\prime}(\mu^{N},\mu).$ (5.29)
Combining estimates (5.28) and (5.29) then yields the desired result.